In the Target Building, the very first casting on the Monolith high bay was performed last week, In total it will be four casting sessions. Last week, we also casted the Admin base slab, the D06 and five slab sections for the D03 Experimental hall. In total we added about 700 m3 of concrete last week.
The beam and head for two cranes arrived to the site last week; the large E01 overheard crane and the E03 slime crane.
In G02 the stub works continue. Installation of conduits, piping and waveguides are being done in 23 of the 27 stubs, connecting the Klystron gallery to the Accelerator tunnel. At the same time the mounting of steel structure in the Klystron gallery continued.
Outside the E-buildings, excavation and piping for water and sewage are ongoing. On Wednesday the last RD pile was installed, making the total number of piles almost 7,000.